# Accurate Chemistry Collection: Coupled cluster atomization energies for broad chemical space

### v1

@@ Methods
The data records for 76,879 molecules of the first public version of MSR-ACC are available in Zenodo (https://doi.org/10.5281/zenodo.15387280) under the CDLA-Permissive-2.0 license. To aid machine-learning applications, we release also training and validation subsets of the full dataset, which constitute canonical 99% and 1% splits, respectively, of MSR-ACC/TAE25 after removing the overlap with popular W4-17 and GMTKN55 benchmark sets. The overlap contains any atomization reaction where the reconstructed molecular graph of the molecules match.
@@
It is well established that the CCSD(T) method cannot generally achieve chemical accuracy for systems dominated by moderate-to-severe nondynamical correlation effects.22, 24, 25, 23, 9, 21, 26, 83, 84, 85, 86, 87 It is therefore essential to estimate the contributions to the TAEs from post-CCSD(T) excitations. A highly successful a priori predictor for such contributions is the so-called %TAE[(T)] diagnostic,22, 23, 9 which is defined as the percentage of the TAE accounted for by parenthetical connected triple excitations:

### v2

@@ Methods
The data records for 76,879 molecules of the first public version of MSR-ACC are available in Zenodo (https://doi.org/10.5281/zenodo.15387280) under the CDLA-Permissive-2.0 license. To aid machine-learning applications,83 we release also training and validation subsets of the full dataset, which constitute canonical 99% and 1% splits, respectively, of MSR-ACC/TAE25 after removing the overlap with popular W4-17 and GMTKN55 benchmark sets. The overlap contains any atomization reaction where the reconstructed molecular graph of the molecules match.
@@
It is well established that the CCSD(T) method cannot generally achieve chemical accuracy for systems dominated by moderate-to-severe nondynamical correlation effects.22, 24, 25, 23, 9, 21, 26, 84, 85, 86, 87, 88 It is therefore essential to estimate the contributions to the TAEs from post-CCSD(T) excitations. A highly successful a priori predictor for such contributions is the so-called %TAE[(T)] diagnostic,22, 23, 9 which is defined as the percentage of the TAE accounted for by parenthetical connected triple excitations: